y = 2x2 – x + 4 at x = 3.
y = 11x – 14
y = 4x + 14
y = 4x + 3
y = 11x + 3

Answer by Alan3354(69443)

y = 2x2 – x + 4 at x = 3
----
Use ^ (Shift 6) for exponents.
y = 2x^2 – x + 4 at x = 3
y = 2*9 - 3 + 4
y = 19
------------
11x - 14 also = 19 at x=3